# Locally Repairable Convertible Codes: Improved Lower Bound and General Construction

### v1

@@ II-B Convertible Codes
Recently, [11] generalizes the framework for studying code conversions proposed by [32], allowing all initial and final codes to have arbitrary parameters and specifying the conversion procedure.
Definition 2 (Generalized Convertible Code [11]):
Let $t_{1},t_{2}$ be two positive integers. A $(t_{1},t_{2})_{q}$ generalized convertible code $\bm{\mathcal{C}}$ over $\mathbb{F}_{q}$ consists of:
1) $t_{1}$ initial codes $\mathcal{C}_{I_{1}},\mathcal{C}_{I_{2}},\cdots,\mathcal{C}_{I_{t_{1}}}$ and $t_{2}$ final codes $\mathcal{C}_{F_{1}},\mathcal{C}_{F_{2}},\cdots,\mathcal{C}_{F_{t_{2}}}$, where $\mathcal{C}_{I_{i}}$ is an $[n_{I_{i}},k_{I_{i}}]_{q}$ code, and $\mathcal{C}_{F_{j}}$ is an $[n_{F_{j}},k_{F_{j}}]_{q}$ code, such that
$\sum_{i\in[t_{1}]}k_{I_{i}}=\sum_{j\in[t_{2}]}k_{F_{j}};$
2) A conversion procedure is determined by a function $\sigma=(\sigma_{1},\sigma_{2},\cdots,\sigma_{t_{2}})$ from $\bigotimes_{i\in[t_{1}]}\mathcal{C}_{I_{i}}\rightarrow\bigotimes_{j\in[t_{2}]}\mathcal{C}_{F_{j}}$, i.e.,
[…]
Specifically, a concrete conversion procedure is realized in three steps.
@@
For the convenience of analysis, we use two-dimensional indices to denote the code symbols, i.e., the codeword $,\bm{c}_{I_{i}}=(c_{i,1},c_{i,2},$$\cdots,c_{i,n_{I_{i}}})\in\mathcal{C}_{I_{i}}$ for $i\in[t_{1}]$.
Step 2: Reading symbols from initial codes. For $j\in[t_{2}]$, the function $\sigma_{j}$ reads some code symbols from each initial code $\mathcal{C}_{I_{i}},1\leqslant i\leqslant t_{1}$, whose indices are denoted by $\mathcal{R}_{1,j},\mathcal{R}_{2,j},\cdots,\mathcal{R}_{t_{1},j}$ where $\mathcal{R}_{i,j}\subseteq\{i\}\times[n_{I_{i}}]$, respectively.
Step 3: Conversion. For any $j\in[t_{2}]$, choose $u_{j}$ unchanged symbols $\mathcal{C}_{I_{1}}|_{\mathcal{U}_{1,j}},\mathcal{C}_{I_{2}}|_{\mathcal{U}_{2,j}},\cdots,\mathcal{C}_{I_{t_{1}}}|_{\mathcal{U}_{t_{1},j}}$. And then apply $\sigma_{j}$ to the read symbols $\mathcal{C}_{I_{1}}|_{\mathcal{R}_{1,j}},\mathcal{C}_{I_{2}}|_{\mathcal{R}_{2,j}},\cdots,\mathcal{C}_{I_{t_{1}}}|_{\mathcal{R}_{t_{1},j}}$ to generate $n_{F_{j}}-u_{j}$ symbols
$\mathcal{C}_{F_{j}}|_{\mathcal{W}_{j}}=\left\{\sigma_{j}(\bm{x}):\bm{x}\in\bigotimes_{i\in[t_{1}]}\mathcal{C}_{I_{i}}|_{\mathcal{R}_{i,j}}\right\},$
@@
$\mathcal{C}_{F_{j}}$ =$=\left\{(\bm{x}_{1},\sigma_{j}(\bm{x}_{2})):\bm{x}\in\bigotimes_{i\in[t_{1}]}\mathcal{C}_{I_{i}},\bm{x}_{1}=\bm{x}|_{\bigcup_{i\in[t_{1}]}\mathcal{U}_{i,j}},\bm{x}_{2}=\bm{x}|_{\bigcup_{i\in[t_{1}]}\mathcal{R}_{i,j}}\right\}$
@@
$\mathcal{F}=\bigcup_{i\in[t_{1}]}\mathcal{U}_{i,j}\cup\mathcal{W}_{j}$ (4)
such that for every $t_{1}$ initial codewords $\bm{c}_{I_{i}}\in\mathcal{C}_{I_{i}}$, $i\in[t_{1}]$, there are $t_{2}$ unique final codewords $\bm{c}_{F_{j}}\in\mathcal{C}_{F_{j}}$ for $j\in[t_{2}]$ satisfying
$\bm{c}_{I_{i}}|_{\mathcal{U}_{i,j}}=\bm{c}_{F_{j}}|_{\mathcal{U}_{i,j}}.$
When initial codes and final codes are both MDS codes, the corresponding convertible code $\bm{\mathcal{C}}$ is called an MDS convertible code. When initial codes and final codes are both LRCs, the corresponding convertible code $\bm{\mathcal{C}}$ is called an locally repairable convertible code (LRCC). In this paper, we mainly consider that all initial codes are with same parameters in the merge regime, i.e., code length $n_{I}$, dimension $k_{I}$ and $(r_{I},\delta_{I})$-locality, which is called a $(t,1)_{q}$ LRCC with parameters $(n_{I},k_{I},r_{I},\delta_{I};n_{F},k_{F},r_{F},\delta_{F})_{a}$ where the subscript $a$ means the initial and final codes are with all symbols locality. If some parameters are unknown, i.e., $r_{I},\delta_{I}$, then we say the above LRCC is with parameters $(n_{I},k_{I},*,*;n_{F},k_{F},r_{F},\delta_{F})_{a}$.
One indicator of the quality of a convertible code is the read and write access costs, which measure the number of symbols affected by the conversion procedure.
@@
For a $(t_{1},t_{2})_{q}$ generalized convertible code $\bm{\mathcal{C}}$, the access cost $\rho$ is defined as the sum of read access cost $\rho_{r}=\sum_{i\in[t_{1}]}|\bigcup_{j\in[t_{2}]}\mathcal{R}_{i,j}|$ and write access cost $\rho_{w}=\sum_{j\in[t_{2}]}(n_{F_{j}}-\sum_{i\in[t_{1}]}|\mathcal{U}_{i,j}|)$, where the read access cost is the number of symbols in initial codes used in the conversion procedure functions and the write access cost counts the number of symbols written in the final codes.
In [23], Kong proposed a lower bound on the access cost of a LRCC with the condition that the final code $\mathcal{C}_{F}$ is an $[n,k,d]_{q}$-LRC with $(r,\delta)_{a}$-locality.
Theorem 1 ([23, Theorem IV.4]):
@@
Write $\Delta=n_{F}-2d-((t-1)k+\lceil\frac{(t-1)k}{r}\rceil)+3$. Then, the read access cost is
$\rho_{r}\geqslant\begin{cases}tk,&\mbox{if }\Delta\leqslant 0\text{ or }d>n_{I}-k+1;\\ t(k-\lceil\frac{r\Delta}{r+1}\rceil),&\mbox{otherwise}.\end{cases}$
@@
When $r=k$, the preceding bound is the same as the lower bound of access cost for MDS convertible codes proposed by [32, Theorem 9], since the $[n,k,d]_{q}$ MDS codes are optimal LRCs with $(k,d)_{a}$-locality.
@@ III Improved Lower Bound on Access Cost
In this section, we improve Theorem 1 by establishing a new lower bound that depends on both $r$ and $\delta$. We first derive an upper bound on the number of unchanged symbols for each initial code.
@@
Let $\bm{\mathcal{C}}$ be a $(t,1)_{q}$ LRCC with parameters $(n_{I},k,*,*;n_{F},tk,r,\delta)_{a}$. Assume that the final code $\mathcal{C}_{F}$ has minimum distance $d$. Then, for every $1\leqslant i\leqslant t$, we have
$|\mathcal{U}_{i,1}|\leqslant n_{F}-d-(t-1)k+1-\left(\left\lceil\frac{(t-1)k}{r}\right\rceil-1\right)(\delta-1).$
@@
Without loss of generality, assume $i=1$. Let $\tilde{\mathcal{C}}$ be the subcode of the final code $\mathcal{C}_{F}$ that is obtained by converting the initial codewords $\bm{c}_{1},\bm{c}_{2},\dots,\bm{c}_{t}$, where $\bm{c}_{1}=\bm{\alpha}$ is a constant vector and $\bm{c}_{i}\in\mathcal{C}_{I_{i}}$ for $i\in[t]\setminus\{1\}$. Let $d(\tilde{\mathcal{C}})$ denote the minimum Hamming distance of $\tilde{\mathcal{C}}$. Then we have $d(\tilde{\mathcal{C}})\geqslant d$ and $|\tilde{\mathcal{C}}|=q^{(t-1)k}$. Define
$N\triangleq\{i\in\mathcal{F}:i\in\mathcal{U}_{1,1}\text{ or exists a set }\mathrm{Loc}(i)\subseteq\mathcal{U}_{1,1}\},$
where $\mathcal{F}$ is defined by (4) and $N^{C}\triangleq\mathcal{F}\setminus N$. Then, for each $i\in N^{C}$ and corresponding $\mathrm{Loc}(i)$, we claim that $\mathrm{Loc}(i)\cap N^{C}\neq\emptyset$. Otherwise, assume that there is a $j\in N^{C}$ such that $\mathrm{Loc}(j)\subseteq N$. Then, each symbol whose index is in $\mathrm{Loc}(j)$ can be recovered by symbols with indices in $\mathcal{U}_{1,1}$, implying that the $j$th symbol can be recovered by symbols with indices in $\mathcal{U}_{1,1}$. This leads to $j\in N$, a contradiction.
@@
On the other hand, since for each $i\in N^{C}$, $\mathrm{Loc}(i)$ satisfies $\mathrm{Loc}(i)\cap N^{C}\neq\emptyset$, and $\mathrm{Loc}(i)\subseteq S_{i}$, where $S_{i}$ is an $(r,\delta)$-repair set for the code symbol $c_{i}$, we have $d(\tilde{\mathcal{C}}|_{S_{i}})\geqslant d(\mathcal{C}_{F}|_{S_{i}})\geqslant\delta$ and $\tilde{\mathcal{C}}|_{S_{i}\cap N}$ is a constant vector. It follows that
$d(\tilde{\mathcal{C}}|_{N^{C}\cap S_{i}})=d(\tilde{\mathcal{C}}|_{S_{i}})\geqslant\delta.$
Hence, $\tilde{\mathcal{C}}|_{N^{C}}$ can be regarded as an $(n_{F}-|N|,(t-1)k)$-LRC with $(r,\delta)_{a}$-locality. By Lemma 2, we obtain
$\begin{split}d(\tilde{\mathcal{C}})&=d(\tilde{\mathcal{C}}|_{N^{C}})\\ &\leqslant(n_{F}-|N|)-(t-1)k+1-\left(\left\lceil\frac{(t-1)k}{r}\right\rceil-1\right)(\delta-1).\end{split}$ (5)
The desirable result follows from $d(\tilde{\mathcal{C}})\geqslant d$, $|\mathcal{U}_{1,1}|\leqslant|N|$ and (5). ∎
Next, we provide a lower bound for the number of read symbols from each initial code.
@@
$\Delta_{i}\triangleq|\mathcal{U}_{i,1}\setminus\mathcal{R}_{i,1}|-d+1.$ (6)
Let $B$ be the largest subset of $\mathcal{U}_{i,1}\setminus\mathcal{R}_{i,1}$ such that
@@
    $B=\cup_{1\leqslant t\leqslant\tau}B_{t}$ with $B_{t_{1}}\cap B_{t_{2}}=\emptyset$, $t_{1}\neq t_{2}$ and $\tau\leqslant\lfloor\frac{\Delta_{i}}{r+\delta-1}\rfloor$;
@@
    For each $1\leqslant t\leqslant\tau$ and $j_{1},j_{2}\in B_{t}$ with $j_{1}\neq j_{2}$, we may set $\mathrm{Loc}(j_{1})$ and $\mathrm{Loc}(j_{2})$ as the same set, i.e., $\mathrm{Loc}(j_{1})=\mathrm{Loc}(j_{2});$
  • (iii)
    For each $1\leqslant t\leqslant\tau$ and $j\in B_{t}$, we have one possible $\mathrm{Loc}(j)$ satisfying
    $\mathrm{Loc}(j)\cap B=\emptyset.$
Then, for each $1\leqslant i\leqslant t$, we have
$|\mathcal{R}_{i,1}|\geqslant\begin{cases}k,&\mbox{if }\Delta_{i}\leqslant 0;\\ k-\Delta_{i}+|B|,&\mbox{otherwise}.\end{cases}$
Moreover, if $d>n_{I}-k+1$, then $\Delta_{i}\leqslant 0$ for $1\leqslant i\leqslant t$ and therefore, $|\mathcal{R}_{i,1}|\geqslant k$.
@@
Without loss of generality, assume $i=1$. Let $\tilde{\mathcal{C}}$ be the subcode of the final code $\mathcal{C}_{F}$ that arises from converting the initial codewords $\bm{c}_{1},\bm{c}_{2},\dots,\bm{c}_{t}$, where $\bm{c}_{i}=\bm{\beta}_{i}$ are constant vectors for $i=2,\dots,t$, and $\bm{c}_{i}\in\mathcal{C}_{I_{i}}$ for $i\in[t]$. Let $d(\tilde{\mathcal{C}})$ denote the minimum distance of $\tilde{\mathcal{C}}$. Then we have $d(\tilde{\mathcal{C}})\geqslant d$ and $|\tilde{\mathcal{C}}|=q^{k}$.
@@
For any $\bm{v}\in\tilde{\mathcal{C}}$, each symbol of $\bm{v}$ is either inherited from $\bm{c}_{1}|_{\mathcal{U}_{1,1}},\bm{\beta}_{2}|_{\mathcal{U}_{2,1}},\dots,\bm{\beta}_{t}|_{\mathcal{U}_{t,1}}$ or determined by symbols from $\bm{c}_{1}|_{\mathcal{R}_{1,1}},\bm{\beta}_{2}|_{\mathcal{R}_{2,1}},\dots,\bm{\beta}_{t}|_{\mathcal{R}_{t,1}}$. Therefore, the non-constant symbols of the codewords in $\tilde{\mathcal{C}}$ are either inherited from $\bm{c}_{1}|_{\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}}$ or determined as functions of symbols in $\bm{c}_{1}|_{\mathcal{R}_{1,1}}$.
@@
According to Lemma 1, we will pick a subset $S\subseteq\mathcal{F}$ of size at least $n_{F}-d+1$ and prove the lower bound by analyzing the relationships among codeword symbols in $\bm{c}_{1}|_{\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}}$, $\bm{c}_{1}|_{\mathcal{R}_{1,1}}$ and $\bm{v}|_{S}$.
@@
Case 1: $\Delta_{1}\leqslant 0$, where $\Delta_{1}$ is defined by (6). In this case, $\mathcal{F}\setminus(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})$ has size at least $n_{F}-d+1\geqslant n_{F}-d(\tilde{\mathcal{C}})+1$ by $|\mathcal{F}|=n_{F}$. Take $S=\mathcal{F}\setminus(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})$, we have $|\tilde{\mathcal{C}}|_{S}|=|\tilde{\mathcal{C}}|=q^{k}$. Note that $\bm{c}_{1}|_{\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}}=\bm{v}|_{\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}}$. Thus, by $S\cap(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})=\emptyset$, non-constant symbols in $\bm{v}|_{S}$ are functions of symbols in $\bm{c}_{1}|_{\mathcal{R}_{1,1}}$. Therefore, $|\tilde{\mathcal{C}}|_{S}|$ is upper bounded by $q^{|\mathcal{R}_{1,1}|}$, which implies that $|\mathcal{R}_{1,1}|\geqslant k$.
Case 2: $\Delta_{1}>0$. Let $N=\bigcup_{j\in B}\mathrm{Loc}(j)$. Since the final code $\mathcal{C}_{F}$ has $(r,\delta)_{a}$-locality, we have

### v2

@@ II-B Convertible Codes
Recently, [13] generalized the framework for studying code conversions proposed by [37], allowing all initial and final codes to be two sets of codes with arbitrary parameters.
Definition 2 (Irregular Conversion Procedure [13]):
Let $t_{1},t_{2}$ be two positive integers. An irregular conversion procedure from $t_{1}$ initial codes $\mathcal{C}_{I_{1}},\mathcal{C}_{I_{2}},\cdots,\mathcal{C}_{I_{t_{1}}}$ to $t_{2}$ final codes $\mathcal{C}_{F_{1}},\mathcal{C}_{F_{2}},\cdots,\mathcal{C}_{F_{t_{2}}}$, where $\mathcal{C}_{I_{i}}$ is an $n_{I_{i}},k_{I_{i}}]_{q}$ code with $r_{I_{i}}\triangleq n_{I_{i}}-k_{I_{i}}$ for $i\in[t_{1}]$, and $\mathcal{C}_{F_{j}}$ is an $n_{F_{j}},k_{F_{j}}]_{q}$ code with $r_{F_{j}}\triangleq n_{F_{j}}-k_{F_{j}}$ for $j\in[t_{2}]$ satisfying
$\sum_{i\in[t_{1}]}k_{I_{i}}=\sum_{j\in[t_{2}]}k_{F_{j}},$ (6)
is a function $\sigma=(\sigma_{1},\sigma_{2},\cdots,\sigma_{t_{2}})$ from $\bigotimes_{i\in[t_{1}]}\mathcal{C}_{I_{i}}\rightarrow\bigotimes_{j\in[t_{2}]}\mathcal{C}_{F_{j}}$, i.e.,
$\sigma_{j}:\bigotimes_{i\in[t_{1}]}\mathcal{C}_{I_{i}}\rightarrow\mathcal{C}_{F_{j}},\quad 1\leqslant j\leqslant t_{2}.$
For the convenience of analysis, we use two-dimensional indices to denote the code symbols, i.e., the codeword $\bm{c}_{I_{i}}=(c_{i,1},c_{i,2},$ $\cdots,c_{i,n_{I_{i}}})\in\mathcal{C}_{I_{i}}$ for $i\in[t_{1}]$. A concrete conversion procedure is realized in three steps as follows:
@@
Step 2: Reading symbols from initial codes. For $j\in[t_{2}]$, the function $\sigma_{j}$ reads some code symbols from each initial code $\mathcal{C}_{I_{i}},1\leqslant i\leqslant t_{1}$, whose indices are denoted by $\mathcal{D}_{1,j},\mathcal{D}_{2,j},\cdots,\mathcal{D}_{t_{1},j}$ where $\mathcal{D}_{i,j}\subseteq\{i\}\times[n_{I_{i}}]$, respectively.
Step 3: Conversion. For any $j\in[t_{2}]$, choose $u_{j}$ unchanged symbols $\mathcal{C}_{I_{1}}|_{\mathcal{U}_{1,j}},\mathcal{C}_{I_{2}}|_{\mathcal{U}_{2,j}},\cdots,\mathcal{C}_{I_{t_{1}}}|_{\mathcal{U}_{t_{1},j}}$. And then apply $\sigma_{j}$ to the read symbols $\mathcal{C}_{I_{1}}|_{\mathcal{D}_{1,j}},\mathcal{C}_{I_{2}}|_{\mathcal{D}_{2,j}},\cdots,\mathcal{C}_{I_{t_{1}}}|_{\mathcal{D}_{t_{1},j}}$ to generate $n_{F_{j}}-u_{j}$ symbols
$\mathcal{C}_{F_{j}}|_{\mathcal{W}_{j}}=\left\{\sigma_{j}(\bm{x}):\bm{x}\in\bigotimes_{i\in[t_{1}]}\mathcal{C}_{I_{i}}|_{\mathcal{D}_{i,j}}\right\},$ (7)
@@
$\mathcal{C}_{F_{j}}$ $=\left\{(\bm{x}_{1},\sigma_{j}(\bm{x}_{2})):\bm{x}\in\bigotimes_{i\in[t_{1}]}\mathcal{C}_{I_{i}},\bm{x}_{1}=\bm{x}|_{\bigcup_{i\in[t_{1}]}\mathcal{U}_{i,j}},\bm{x}_{2}=\bm{x}|_{\bigcup_{i\in[t_{1}]}\mathcal{D}_{i,j}}\right\}$ (8)
$\subseteq\bigotimes_{i\in[t_{1}]}\mathcal{C}_{I_{i}}|_{\mathcal{U}_{i,j}}\otimes\mathcal{C}_{F_{j}}|_{\mathcal{W}_{j}}$
@@
$\bigcup_{i\in[t_{1}]}\mathcal{U}_{i,j}\cup\mathcal{W}_{j}$ (9)
such that
$\mathcal{C}_{I_{i}}|_{\mathcal{U}_{i,j}}=\mathcal{C}_{F_{j}}|_{\mathcal{U}_{i,j}}.$ (10)
Definition 3 (Irregular Convertible Code [13]):
Let $t_{1},t_{2}$ be two positive integers. A $(t_{1},t_{2})_{q}$ irregular convertible code $\bm{\mathcal{C}}$ over $\mathbb{F}_{q}$ consists of: (1) $t_{1}$ initial codes $\mathcal{C}_{I_{1}},\mathcal{C}_{I_{2}},\cdots,\mathcal{C}_{I_{t_{1}}}$ and $t_{2}$ final codes $\mathcal{C}_{F_{1}},\mathcal{C}_{F_{2}},\cdots,\mathcal{C}_{F_{t_{2}}}$, where $\mathcal{C}_{I_{i}}$ is an $n_{I_{i}},k_{I_{i}}]_{q}$ code for $i\in[t_{1}]$, and $\mathcal{C}_{F_{j}}$ is an $n_{F_{j}},k_{F_{j}}]_{q}$ code for $j\in[t_{2}]$ satisfying $\sum_{i\in[t_{1}]}k_{I_{i}}=\sum_{j\in[t_{2}]}k_{F_{j}}$; (2) an irregular conversion procedure $\sigma$ defined by Definition 2.
When both the initial and final codes are MDS codes, the corresponding convertible code $\bm{\mathcal{C}}$ is called an MDS convertible code. When both the initial and final codes are LRCs, the corresponding convertible code $\bm{\mathcal{C}}$ is called an locally repairable convertible code (LRCC). In this paper, we mainly consider the case where all initial codes have the same parameters in the merge regime, i.e., code length $n_{I}$, dimension $k_{I}$ and $(r_{I},\delta_{I})$-locality, which is called a $(t,1)_{q}$ LRCC with parameters $(n_{I},k_{I},r_{I},\delta_{I};n_{F},k_{F},r_{F},\delta_{F})_{a}$ where the subscript $a$ means both the initial and final codes are with all symbols locality. If some parameters $r_{I},\delta_{I}$ are unknown, then we say the above LRCC is with parameters $(n_{I},k_{I},*,*;n_{F},k_{F},r_{F},\delta_{F})_{a}$.
One merit factor of a convertible code is the read and write access costs, which measure the number of symbols accessed by the conversion procedure.
@@
For a $(t_{1},t_{2})_{q}$ irregular convertible code $\bm{\mathcal{C}}$, the access cost $\rho$ is defined as the sum of read access cost $\rho_{r}=\sum_{i\in[t_{1}]}|\bigcup_{j\in[t_{2}]}\mathcal{D}_{i,j}|$ and write access cost $\rho_{w}=\sum_{j\in[t_{2}]}(n_{F_{j}}-\sum_{i\in[t_{1}]}|\mathcal{U}_{i,j}|)$, where the read access cost is the number of symbols in initial codes used in the conversion procedure functions and the write access cost counts the number of symbols written in the final codes.
In [27], Kong proposed a lower bound on the access cost of an LRCC with the condition that the final code $\mathcal{C}_{F}$ is an $n,k,d]_{q}-$LRC with $(r,\delta)_{a}$-locality.
Theorem 1 ([27, Theorem IV.4]):
@@
Write $\overline{\Delta}=n_{F}-2d-((t-1)k+\lceil\frac{(t-1)k}{r}\rceil)+3$. Then, the read access cost is
$\rho_{r}\geqslant\begin{cases}tk,&\mbox{if }\overline{\Delta}\leqslant 0\text{ or }d>n_{I}-k+1{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}\\ t(k-\lceil\frac{r\overline{\Delta}}{r+1}\rceil),&\mbox{otherwise}.\end{cases}$
@@
When $r=k$, the preceding bound is equivalent to the lower bound on the access cost for MDS convertible codes proposed by [37, Theorem 9], since the $n,k,d]_{q}$ MDS codes are optimal LRCs with $(k,d)_{a}$-locality.
@@ III Improved Lower Bound on Access Cost
In this section, we improve Theorem 1 by establishing a new lower bound that depends on both $r$ and $\delta$. At first, we derive an upper bound on the number of unchanged symbols for each initial code.
@@
Let $\bm{\mathcal{C}}$ be a $(t,1)_{q}$ LRCC with parameters $(n_{I},k,*,*;n_{F},tk,r,\delta)_{a}$. Assume that the final code $\mathcal{C}_{F}$ has the minimum distance $d$ and the optimal Hamming distance $d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)=n-k+1-\phi(k,r,\delta)$. Then, for every $1\leqslant i\leqslant t$,
$|\mathcal{U}_{i,1}|\leqslant n_{F}-d-(t-1)k+1-\phi((t-1)k,r,\delta).$
@@
[…]
$N\triangleq\{i\in\mathcal{F}_{1}:i\in\mathcal{U}_{1,1}\text{ or exist a set }\mathrm{Loc}(i)\subseteq\mathcal{U}_{1,1}\},$
where $\mathcal{F}_{1}$ is defined by (9).
@@
On the other hand, for each $i\in N^{C}\triangleq\mathcal{F}_{1}\setminus N$ and its corresponding $\mathrm{Loc}(i)$, we claim that $\mathrm{Loc}(i)\cap N^{C}\neq\emptyset$. Otherwise, assume that there is a $j\in N^{C}$ such that $\mathrm{Loc}(j)\subseteq N$. Then, each symbol whose index is in $\mathrm{Loc}(j)$ can be recovered by symbols with indices in $\mathcal{U}_{1,1}$, implying that the $j$th symbol can be recovered by symbols with indices in $\mathcal{U}_{1,1}$. This leads to $j\in N$, a contradiction.
Since for each $i\in N^{C}$, $\mathrm{Loc}(i)$ satisfies $\mathrm{Loc}(i)\cap N^{C}\neq\emptyset$, and $\mathrm{Loc}(i)\subseteq S_{i}$, where $S_{i}$ is an $(r,\delta)$-repair set for the code symbol $c_{i}$, we have $d(\tilde{\mathcal{C}}|_{S_{i}})\geqslant d(\mathcal{C}_{F}|_{S_{i}})\geqslant\delta$ and $\tilde{\mathcal{C}}|_{S_{i}\cap N}$ is a constant vector. It follows that d(~C—_N^C∩S_i) = d(~C—_S_i) ⩾δ. Hence, $\tilde{\mathcal{C}}|_{N^{C}}$ can be regarded as an $(n_{F}-|N|,(t-1)k)$-LRC with $(r,\delta)_{a}$-locality. We obtain
$\begin{split}d(\tilde{\mathcal{C}})&=d(\tilde{\mathcal{C}}|_{N^{C}})\\ &\leqslant d_{\mathrm{opt}}^{(q)}(n_{F}-|N|,(t-1)k,r,\delta)\\ &\leqslant(n_{F}-|N|)-(t-1)k+1-\phi((t-1)k,r,\delta).\end{split}$ (11)
The desirable result holds from $d(\tilde{\mathcal{C}})\geqslant d$, $|\mathcal{U}_{1,1}|\leqslant|N|$ and (11). ∎
Secondly, we provide a lower bound on the number of read symbols from each initial code.
@@
$\Delta_{i}\triangleq|\mathcal{U}_{i,1}\setminus\mathcal{D}_{i,1}|-d+1.$ (12)
Let $B=\cup_{1\leqslant t\leqslant\tau}B_{t}$ be the subset of $\mathcal{U}_{i,1}\setminus\mathcal{D}_{i,1}$ with size at most $(\delta-1)\lfloor\frac{\Delta_{i}}{r+\delta-1}\rfloor$ for a positive integer $\tau$, such that
@@
    For each $1\leqslant t\leqslant\tau$, the symbols $c_{i}$ for all $i\in B_{t}$ are in the same repair set $S_{t}$ of the final code $\mathcal{C}_{F}$, and $B_{t_{1}}\cap S_{t_{2}}=\emptyset$ for all $1\leqslant t_{2}<t_{1}\leqslant\tau$;
@@
    For each $1\leqslant t\leqslant\tau$, $|B_{t}|\leqslant\delta-1$, and
    $\left|\left(S_{t}\setminus\bigcup_{1\leqslant j\leqslant t-1}S_{j}\right)\cap(\mathcal{U}_{i,1}\setminus\mathcal{D}_{i,1})\right|<\delta-1,$ (13)
    if $|B_{t}|<\delta-1$.
Then, for each $1\leqslant i\leqslant t$,
$|\mathcal{D}_{i,1}|\geqslant\begin{cases}k,&\mbox{if }\Delta_{i}\leqslant 0{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}\\ k-\Delta_{i}+|B|,&\mbox{otherwise}.\end{cases}$
Moreover, if $d>n_{I}-k+1$, then $\Delta_{i}\leqslant 0$ for $1\leqslant i\leqslant t$ and therefore, $|\mathcal{D}_{i,1}|\geqslant k$.
@@
[…]
@@
For any $\bm{v}\in\tilde{\mathcal{C}}$, each symbol of $\bm{v}$ is either inherited from $\bm{c}_{1}|_{\mathcal{U}_{1,1}},\bm{\beta}_{2}|_{\mathcal{U}_{2,1}},\dots,\bm{\beta}_{t}|_{\mathcal{U}_{t,1}}$ or determined by symbols from $\bm{c}_{1}|_{\mathcal{D}_{1,1}},\bm{\beta}_{2}|_{\mathcal{D}_{2,1}},\dots,\bm{\beta}_{t}|_{\mathcal{D}_{t,1}}$. Therefore, the non-constant symbols of the codewords in $\tilde{\mathcal{C}}$ are either inherited from $\bm{c}_{1}|_{\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1}}$ or are functions of symbols in $\bm{c}_{1}|_{\mathcal{D}_{1,1}}$.
@@
According to Lemma 1, we will pick a subset $M\subseteq\mathcal{F}_{1}$ of size at least $n_{F}-d+1$ and prove the lower bound by analyzing the relationships among codeword symbols in $\bm{c}_{1}|_{\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1}}$, $\bm{c}_{1}|_{\mathcal{D}_{1,1}}$ and $\bm{v}|_{M}$.
@@
Case 1: $\Delta_{1}\leqslant 0$, i.e., $|\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1}|\leqslant d-1$ by (12). In this case, $\mathcal{F}_{1}\setminus(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1})$ has size at least $n_{F}-d+1\geqslant n_{F}-d(\tilde{\mathcal{C}})+1$ by $|\mathcal{F}_{1}|=n_{F}$. Take $M=\mathcal{F}_{1}\setminus(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1})$, we have $|\tilde{\mathcal{C}}|_{M}|=|\tilde{\mathcal{C}}|=q^{k}$. Note that $\bm{c}_{1}|_{\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1}}=\bm{v}|_{\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1}}$. Thus, by $M\cap(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1})=\emptyset$, non-constant symbols in $\bm{v}|_{M}$ are functions of symbols in $\bm{c}_{1}|_{\mathcal{D}_{1,1}}$. Therefore, $|\tilde{\mathcal{C}}|_{M}|$ is upper bounded by $q^{|\mathcal{D}_{1,1}|}$, which implies that $|\mathcal{D}_{1,1}|\geqslant k$.
Case 2: $\Delta_{1}>0$, i.e., $|\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1}|>d-1$ by (12). By condition (ii), $|B_{i}|\leqslant\delta-1$. Denote $\mathcal{A}_{1}\triangleq\{i:|B_{i}|=\delta-1,i\in[\tau]\}$ and $\mathcal{A}_{2}\triangleq\{i:|B_{i}|<\delta-1,i\in[\tau]\}$. Thus, $B=\cup_{i\in\mathcal{A}_{1}\cup\mathcal{A}_{2}}B_{i}$. Let N = ⋃_ i∈A_1 (S_i ∖B_i). Thus,